# Robust methods for Darcy–Forchheimer equationsDas, Hutridurga, Pani, Ruiz-Baier

### v1

@@ 6 Numerical results
The purpose of this section is to experimentally validate the theoretical results presented in the previous sections. We verify that the convergence rates are robust in the proposed finite element method, and also illustrate the performance of proposed schemes in typical porous media flow problems in 2D and 3D. The numerical implementation is based on the open-source finite element framework Gridap [4].
@@
144 0.7071 8.18e-02 $\star$ 2.34e-01 $\star$ 1.49e-14 3
400 0.3536 3.11e-02 1.395 5.90e-02 1.991 1.26e-14 3
1296 0.1768 9.89e-03 1.654 1.13e-02 2.387 3.69e-14 4
4624 0.0884 2.81e-03 1.818 1.95e-03 2.530 6.67e-14 4
17424 0.0442 7.48e-04 1.908 3.68e-04 2.409 1.67e-13 4
67600 0.0221 1.93e-04 1.954 8.09e-05 2.184 3.84e-13 4
266256 0.0110 4.90e-05 1.977 1.95e-05 2.052 8.43e-13 4
Table 6.1: Example 1. Error history against smooth manufactured solutions (errors on a sequence of successively refined grids, convergence rates, and norm of the divergence of the discrete velocity) for different polynomial degrees, and iteration count for the nonlinear Newton–Raphson solver. Here we are just using unity parameters (and $r=3.5$) and the $\mathbf{H}^{3}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and $\mathrm{L}^{2}(\Omega)$ norms for velocity and pressure, respectively.
Figure 6.1: Example 1. Streamlines of the approximate velocity (left) and pressure distribution (right) for the accuracy test of the Darcy–Forchheimer equations with unit parameters.
Example 1 (convergence against smooth solutions in 2D). Consider the unit square domain $\Omega=(0,1)^{2}$ along with the following manufactured solutions
$p(x,y)=\sin(\pi x)\cos(\pi y),\quad\text{ and }\boldsymbol{u}(x,y):=\begin{pmatrix}\cos(\pi x)\sin(\pi y)\\ -\sin(\pi x)\cos(\pi y)\end{pmatrix}$.
We set the left and bottom sides of the square as $\Gamma_{\boldsymbol{u}}$ and the top and right sides as $\Gamma_{p}$. The forcing vector $\boldsymbol{f}$ and the fluid source $g$, as well as the (non-homogeneous) essential boundary condition on the flux, are computed based on the manufactured solutions above. We take, for this first set of examples, simply dimensional unity parameters $\kappa=\mathtt{F}=1$ and index $r=3.5$. The error history associated with the proposed mixed finite element method on a sequence of successively refined partitions of the domain, are presented in Table 6.1, where we also tabulate rates of error decay computed as $\texttt{rate}=\log(e_{(\bullet)}/\tilde{e}_{(\bullet)})[\log(h/\tilde{h})]^{-1}$, where $e,\tilde{e}$ denote errors generated on two consecutive meshes of sizes $h$ and $\tilde{h}$, respectively. The results indicate optimal convergence in all fields and for the two tested polynomial degrees. Moreover, a maximum of four iterations are needed by the Newton–Raphson method to reach a tolerance (either absolute or relative) of $10^{-8}$ on the residual. Sample approximate solutions for velocity and pressure obtained with the second-order method (with $k=1$) are plotted in Figure 6.1. For these results, we consider the norms $\mathbf{H}^{3}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and $\mathrm{L}^{2}(\Omega)$ for velocity and pressure (with no parameter weighting).
DoF $h$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{3,\mathop{\mathrm{div}}\nolimits,\Omega}\!$ rate $\|p-p_{h}\|_{0,\Omega}$ rate $\|P_{h}(\mathop{\mathrm{div}}\nolimits\boldsymbol{u}_{h}-g)\|_{\ell^{\infty}}$
@@
Table 6.2: Example 2. Error history against a fine-mesh reference solution (experimental errors on a sequence of successively refined grids, convergence rates, and norm of the divergence of the discrete velocity) for different polynomial degrees. Here we use $\kappa=10^{-8}$, $\mathtt{F}=10^{4}$ (and $r=3.5$), and the non-weighted (top rows) and parameter-weighted (bottom rows) norms for velocity and pressure.
@@
Example 2 (experimental self-convergence with extreme parameters). For a second run of simulations we use the domain $\Omega=(0,2)\times(0,1)$ without a known analytical solution, setting the problem data as $\boldsymbol{f}=\boldsymbol{0}$, $g=0$, mixed boundary conditions as follows $\boldsymbol{u}\cdot\boldsymbol{n}=2.5y(1-y)$ on the left segment, $\boldsymbol{u}\cdot\boldsymbol{n}=0$ on the top and bottom segments, and $p=0$ on the right edge. The permeability is heterogeneous $\kappa(\boldsymbol{x})=\kappa_{0}(1+\exp(-\frac{1}{2}[10y-5-\sin(10x)]^{2})$, and we consider the more extreme parameter values $\kappa_{0}=10^{-8}$, $\mathtt{F}=10^{4}$, and $r=3.5$. In Table 6.2 we show the error decay where we compute experimental errors against a fine-mesh reference solution $\boldsymbol{u}_{\mathrm{ref}},p_{\mathrm{ref}}$(with two more refinement steps than the finest level). In the first part of the table we use the non-weighted norms and we see very large errors in the pressure, a sub-optimal error decay for the velocity in the second-order case. In contrast, for the second part of the table the errors are now measured in the weighted norms associated with $\mathbf{V}$ and $\widehat{\mathrm{Q}}(\mathcal{T}_{h})$ defined in (4.1) and (4.3), respectively, where we move the scaling of the now variable permeability inside the norm definition. We recall from [5] that in the space $\mathrm{Q}_{h}$ the realisation of the required discrete weighted Laplacian operator that is the Riesz map for $\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{h,\Gamma_{p}}(\Omega)$ is as follows
@@
$\langle-\mathop{\mathrm{div}}\nolimits(\kappa(\boldsymbol{x})\nabla p_{h}),q_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}(\kappa(\boldsymbol{x})\nabla p_{h},\nabla q_{h})_{K}+\sum_{F\in\mathcal{E}_{h}\cup\mathcal{E}_{h}^{\Gamma_{p}}}\frac{1}{h_{F}}(\kappa(\boldsymbol{x})[\![p_{h}]\!],[\![q_{h}]\!])_{F},$
@@
where $\mathcal{E}_{h}$ denotes the set of internal facets, $\mathcal{E}_{h}^{\Gamma_{p}}$ denotes the set of facets lying on the sub-boundary $\Gamma_{p}$, and $\![\bullet$]\!] represents the jump operator across a facet (and adopt the convention that it reduces to the identity if the face lies on the boundary). In turn, for the space $\mathtt{F}^{-\frac{1}{3}}\mathrm{W}^{1,\frac{3}{2}}_{h,\Gamma_{p}}(\Omega)$ the natural operator (duality map) is the nonlinear discrete $s$-Laplacian
@@
$\langle-\mathop{\mathrm{div}}\nolimits(\mathtt{F}^{-1}|\nabla p_{h}|^{s-2}\nabla p_{h}),q_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}\mathtt{F}^{-1}(|\nabla p_{h}|^{s-2}\nabla p_{h},\nabla q_{h})_{K}$
$\quad+\sum_{F\in\mathcal{E}_{h}\cup\mathcal{E}_{h}^{\Gamma_{p}}}\mathtt{F}^{-1}h_{F}^{1-s}(|[\![p_{h}]\!]|^{s-2}[\![p_{h}]\!],[\![q_{h}]\!])_{F},$
@@
with $s=\frac{3}{2}$ (see, e.g., [9]). However, since we require a linear operator to define the preconditioner-induced norm, we use instead the following linearisation around the reference pressure
$\sum_{K\in\mathcal{T}_{h}}\mathtt{F}^{-1}(|\nabla p_{\mathrm{ref}}|^{s-2}\nabla p_{h},\nabla q_{h})_{K}+\sum_{F\in\mathcal{E}_{h}\cup\mathcal{E}_{h}^{\Gamma_{p}}}\mathtt{F}^{-1}(s-1)h_{F}^{1-s}(|[\![p_{\mathrm{ref}}]\!]|^{s-2}[\![p_{h}]\!],[\![q_{h}]\!])_{F}$
$\quad+\sum_{K\in\mathcal{T}_{h}}\mathtt{F}^{-1}(s-2)(|\nabla p_{\mathrm{ref}}|^{s-4}(\nabla p_{\mathrm{ref}}\cdot\nabla p_{h}),(\nabla p_{\mathrm{ref}}\cdot\nabla q_{h}))_{K}.$
@@
is 0.245 (the rule suggested in [35], see also [15], specifies that it has to be larger than 0.1) when taking for $|\boldsymbol{u}|$ its maximum value over $\Omega$ (not known a priori but computed to be 0.49). We depict in Figure 6.3 the approximate solutions, indicating the expected filtration patterns in porous media with large obstacles. The Newton–Raphson solver took eight iterations to converge.
@@
Example 4 (operator preconditioning for the linearised problem). To conclude we conduct a series of computations where the permeability $\kappa$ and the Forchheimer coefficient $r$ are varied over several orders of magnitude, and the Forchheimer index $\mathtt{F}$ is also varied. This is done again using the domain $\Omega=(0,1)^{2}$, the lowest-order finite element family, and mixed boundary conditions as in Example 1. The performance of the numerical method is assessed by comparing an estimate of the condition number of the linearised system matrix (5.3) preconditioned with the weighted duality map (5.4). Table 6.3 illustrates the robustness of the proposed preconditioner across the parametric space. Additionally, the performance of the preconditioner corresponding to the weighted Riesz map (5.8) has also been illustrated in Table 6.4, which seems to perform slightly better than (5.4). However, we remark that, quite similarly to [5], the map (5.4) will be preferable in a multiphysics context where the scaling of the divergence component of the velocity block is not convenient. Another crucial point worth mentioning is that, although extreme parametric values of both $\kappa$ and $\mathtt{F}$ are considered, and they have the opposite limiting directions, with more refinement of the mesh, effect of $\kappa$ seems to dominate the effects of $\mathtt{F}$ in the sense of higher condition numbers.
@@
$2^{-2}$ $2^{-3}$ $2^{-4}$
5 $10^{-8}$ 1 6.61 7.52 8.01
$10^{3}$ 6.61 7.52 8.01
$10^{9}$ 6.61 7.52 8.01
$10^{-4}$ 1 6.41 6.77 6.78
$10^{3}$ 6.41 6.77 6.78
$10^{9}$ 6.41 6.77 6.78
1 1 2.08 2.11 2.13
$10^{3}$ 2.08 2.11 2.13
$10^{9}$ 2.08 2.11 2.13
Table 6.3: Example 4. Estimated condition number of the preconditioned system matrix of the linearised Darcy–Forchheimer equations using the preconditioner induced by (5.4).
@@
$2^{-2}$ $2^{-3}$ $2^{-4}$
5 $10^{-8}$ 1 2.63 2.63 2.63
$10^{3}$ 2.63 2.63 2.63
$10^{9}$ 2.63 2.63 2.63
$10^{-4}$ 1 2.63 2.63 2.63
$10^{3}$ 2.63 2.63 2.63
$10^{9}$ 2.63 2.63 2.63
1 1 2.63 2.63 2.63
$10^{3}$ 2.63 2.63 2.63
$10^{9}$ 2.63 2.63 2.63
Table 6.4: Example 4. Estimated condition number of the preconditioned system matrix of the linearised Darcy–Forchheimer equations using the preconditioner induced by (5.8).

### v2

@@ 6 Numerical results
The purpose of this section is to experimentally validate the theoretical results presented in the previous sections. We verify that the convergence rates are robust in the proposed finite element method in 2D and 3D, and also illustrate the performance of proposed schemes in typical porous media flow problems. The numerical implementation is based on the open-source finite element framework Gridap [4], and the main routines can be publicly accessed from https://github.com/ruizbaier/GridapRobustDarcyForchheimer.jl.
@@
624 0.6124 3.53e-01 $\star$ 6.32e-02 $\star$ 1.48e-14 5
4992 0.3062 9.86e-02 1.840 1.73e-02 1.871 3.33e-14 6
39936 0.1531 2.55e-02 1.950 4.42e-03 1.968 9.14e-14 6
319488 0.0765 6.45e-03 1.986 1.11e-03 1.992 2.25e-13 6
2555904 0.0383 1.62e-03 1.993 2.78e-04 1.996 1.01e-13 7
20447232 0.0191 4.07e-04 1.997 6.95e-05 1.998 2.01e-13 7
Table 6.1: Example 1. Error history against smooth manufactured solutions in 3D (errors on a sequence of successively refined grids, convergence rates, and norm of the divergence of the discrete velocity) for different polynomial degrees, and iteration count for the nonlinear Newton–Raphson solver. Here we use unity parameters (and $r=3$) and the $\mathbf{H}^{3}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and $\mathrm{L}^{2}(\Omega)$ norms for velocity and pressure, respectively.
Figure 6.1: Example 1. Streamlines of the approximate velocity (left) and iso-contours of pressure distribution (right) for the accuracy test of the Darcy–Forchheimer equations in 3D with unit parameters.
Example 1 (convergence against smooth solutions in 3D). Consider the unit cube domain $\Omega=(0,1)^{3}$ along with the following manufactured solutions
$p(x,y,z):=\sin(\pi x)\cos(\pi y)\sin(\pi z),\quad\boldsymbol{u}(x,y,z):=\begin{pmatrix}\cos(\pi x)\sin(\pi y)\sin(\pi z)\\ -\sin(\pi x)\cos(\pi y)\sin(\pi z)\\ \sin(\pi x)\sin(\pi y)\cos(\pi z)\end{pmatrix}$.
We set the left, bottom, and front sides of the cube as $\Gamma_{\boldsymbol{u}}$ and the right, top, and back sides as $\Gamma_{p}$. The forcing vector $\boldsymbol{f}$ and the fluid source $g$, as well as the (non-homogeneous) essential boundary condition on the flux, are computed based on the manufactured solutions above. We take, for this first set of examples, simply dimensional unity parameters $\kappa=\mathtt{F}=1$ and index $r=3$. The error history associated with the proposed mixed finite element method on a sequence of successively refined partitions of the domain, are presented in Table 6.1, where we also tabulate rates of error decay computed as $\texttt{rate}=\log(e_{(\bullet)}/\tilde{e}_{(\bullet)})[\log(h/\tilde{h})]^{-1}$, where $e,\tilde{e}$ denote errors generated on two consecutive meshes of sizes $h$ and $\tilde{h}$, respectively. The results indicate optimal convergence in all fields and for the two tested polynomial degrees. Moreover, for this problem the number of Newton–Raphson iterations required to reach a tolerance (either absolute or relative) of $10^{-8}$ is up to seven. This depends on the initial guess, and here we have used a constant velocity and pressure such that every entry of the vector of degrees of freedom is $10^{-4}$. Sample approximate solutions for velocity and pressure obtained with the second-order method (with $k=1$) are plotted in Figure 6.1. For these results, we consider the norms $\mathbf{H}^{3}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and $\mathrm{L}^{2}(\Omega)$ for velocity and pressure (with no parameter weighting).
@@
Table 6.2: Example 2. Error history against a fine-mesh reference solution (experimental errors on a sequence of successively refined grids, convergence rates, norm of the divergence of the discrete velocity, and Newton iteration count) for different polynomial degrees. Here we use $\kappa=10^{-8}$, $\mathtt{F}=10^{4}$ (and $r=3$), and the non-weighted (top rows) and parameter-weighted (bottom rows) norms for velocity and pressure.
@@
Example 2 (experimental self-convergence with extreme parameters). For a second run of simulations we use the domain $\Omega=(0,2)\times(0,1)$ without a known analytical solution, setting the problem data as $\boldsymbol{f}=\boldsymbol{0}$, $g=0$, mixed boundary conditions as follows $\boldsymbol{u}\cdot\boldsymbol{n}=2.5y(1-y)$ on the left segment, $\boldsymbol{u}\cdot\boldsymbol{n}=0$ on the top and bottom segments, and $p=0$ on the right edge. The permeability is heterogeneous $\kappa(\boldsymbol{x})=\kappa_{0}(1+\exp(-\frac{1}{2}[10y-5-\sin(10x)]^{2})$, and we consider the more pronounced parameter values $\kappa_{0}=10^{-8}$, $\mathtt{F}=10^{4}$, and $r=3$. In Table 6.2 we show the error decay where we compute experimental errors against a fine-mesh reference solution $\boldsymbol{u}_{\mathrm{ref}},p_{\mathrm{ref}}$(with two more refinement steps than the finest level). In the first part of the table we use the non-weighted norms and we see very large errors in the pressure and a sub-optimal error decay for the velocity in the second-order case. In contrast, for the second part of the table the errors are now measured in the weighted norms associated with $\mathbf{V}$ and $\widehat{\mathrm{Q}}(\mathcal{T}_{h})$ defined in (4.1) and (4.3), respectively, where we move the scaling of the now variable permeability inside the norm definition. We recall from [5] that in the space $\mathrm{Q}_{h}$ the realisation of the required discrete weighted Laplacian operator that is the Riesz map for $\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{h,\Gamma_{p}}(\Omega)$ is as follows
@@
$\langle-\mathop{\mathrm{div}}\nolimits(\kappa(\boldsymbol{x})\nabla p_{h}),q_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}(\kappa(\boldsymbol{x})\nabla p_{h},\nabla q_{h})_{K}+\sum_{F\in\mathcal{E}_{h}\cup\mathcal{E}_{h}^{\Gamma_{p}}}\frac{1}{h_{F}}(\kappa(\boldsymbol{x})[\![p_{h}]\!],[\![q_{h}]\!])_{F}.$
@@
Here $\mathcal{E}_{h}$ denotes the set of internal facets, $\mathcal{E}_{h}^{\Gamma_{p}}$ denotes the set of facets lying on the sub-boundary $\Gamma_{p}$, and $\![\bullet$]\!] represents the jump operator across a facet (and adopt the convention that it reduces to the identity if the face lies on the boundary). In turn, for the space $\mathtt{F}^{-\frac{1}{3}}\mathrm{W}^{1,\frac{3}{2}}_{h,\Gamma_{p}}(\Omega)$ the natural operator (duality map) is the nonlinear discrete $r^{\prime}$-Laplacian
@@
$\langle-\mathop{\mathrm{div}}\nolimits(\mathtt{F}^{r-4}|\nabla p_{h}|^{r^{\prime}-2}\nabla p_{h}),q_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}\mathtt{F}^{r-4}(|\nabla p_{h}|^{r^{\prime}-2}\nabla p_{h},\nabla q_{h})_{K}$
$\quad+\sum_{F\in\mathcal{E}_{h}\cup\mathcal{E}_{h}^{\Gamma_{p}}}\!\!\!\mathtt{F}^{r-4}h_{F}^{1-r^{\prime}}(|[\![p_{h}]\!]|^{r^{\prime}-2}[\![p_{h}]\!],[\![q_{h}]\!])_{F},$
@@
with $r^{\prime}=\frac{3}{2}$ (see, e.g., [10]). However, since we require a linear operator to define the preconditioner-induced norm, we use instead the following linearisation around the reference pressure
$\sum_{K\in\mathcal{T}_{h}}\mathtt{F}^{r-4}(|\nabla p_{\mathrm{ref}}|^{r^{\prime}-2}\nabla p_{h},\nabla q_{h})_{K}+\sum_{F\in\mathcal{E}_{h}\cup\mathcal{E}_{h}^{\Gamma_{p}}}\!\!\!\!\!\mathtt{F}^{r-4}(r^{\prime}-1)h_{F}^{1-r^{\prime}}(|[\![p_{\mathrm{ref}}]\!]|^{r^{\prime}-2}[\![p_{h}]\!],[\![q_{h}]\!])_{F}$
$\quad+\sum_{K\in\mathcal{T}_{h}}\mathtt{F}^{r-4}(r^{\prime}-2)(|\nabla p_{\mathrm{ref}}|^{r^{\prime}-4}(\nabla p_{\mathrm{ref}}\cdot\nabla p_{h}),(\nabla p_{\mathrm{ref}}\cdot\nabla q_{h}))_{K}.$
@@
is 0.245 (the rule suggested in [36], see also [16], specifies that it has to be larger than 0.1) when taking for $|\boldsymbol{u}|$ its maximum value over $\Omega$ (not known a priori but computed to be 0.49). We depict in Figure 6.3 the approximate solutions, indicating the expected filtration patterns in porous media with large obstacles. For this set of simulations the Newton–Raphson solver took nine iterations to converge, and we stress that the iteration count is dependent on the initial guess and parameter configuration.
@@
Example 4 (operator preconditioning for the linearised problem). To conclude we conduct a series of computations where the permeability $\kappa$ is varied over several orders of magnitude, as well as the Forchheimer coefficient $\mathtt{F}$. We also take three values of the index $r$. This is done again using the domain $\Omega=(0,1)^{2}$, the lowest-order finite element family, and mixed boundary conditions as in Example 1. The performance of the numerical method is assessed by comparing an estimate of the condition number of the linearised system matrix (5.3) preconditioned with the weighted duality map (5.4), and also inspect the number of MinRes iterations required to solve the preconditioned system. Table 6.3 illustrates the robustness of the proposed preconditioner across the parametric space. Additionally, the performance of the preconditioner corresponding to the weighted Riesz map (5.8) has also been illustrated in Table 6.4, which seems to perform slightly better than (5.4). This is a behaviour very similar to the one reported in [5], but the map (5.4) is preferable in a multiphysics context where the scaling of the divergence component of the velocity block is not feasible. Note also that both preconditioners are perfectly robust on $r$ (the three sub-tables on each table look almost identical in both condition number and MinRes iteration count). They are also robust on $h$ (the growth in the rows is minimal). However, for the preconditioner induced by (5.4), there are some effects of small values of $\kappa$, which are more noticeable in the MinRes iterations than in the condition number. It is evident from Table 6.4 that large values of $\mathtt{F}$ only seem to affect the number of iterations when $\kappa$ is large too, but in applications one typically expects large $\mathtt{F}$ associated with small $\kappa$.
@@
$\sqrt{2}^{-2}$ $\sqrt{2}^{-3}$ $\sqrt{2}^{-4}$ $\int^{\int^{X}}$
4 $10^{-9}$ $\int^{X}$ $10^{-9}$ 3.74 (23) 4.18 (31) 4.41 (33)
1 3.74 (23) 4.18 (31) 4.41 (33)
$10^{4}$ 3.74 (23) 4.18 (31) 4.41 (33)
$10^{-4}$ $10^{-9}$ $\int^{X}$ 3.65 (25) 3.82 (29) 3.57 (30)
1 3.65 (25) 3.82 (29) 3.57 (30)
$10^{4}$ 3.66 (25) 3.84 (29) 3.57 (30)
1 $10^{-9}$ $\int^{X}$ 1.23 (11) 1.21 (11) 1.21 (11)
1 1.23 (11) 1.22 (11) 1.22 (11)
$10^{4}$ 3.84 (27) 3.16 (26) 3.05 (26)
Table 6.3: Example 4. Estimated condition number and MinRes iterations required to converge, associated with the preconditioned system matrix of the linearised Darcy–Forchheimer equations using the preconditioner induced by (5.4).
@@
$\sqrt{2}^{-2}$ $\sqrt{2}^{-3}$ $\sqrt{2}^{-4}$ $\int^{\int^{X}}$
4 $10^{-9}$ $\int^{X}$ $10^{-9}$ 1.20 (5) 1.20 (1) 1.20 (1)
1 1.20 (5) 1.20 (1) 1.20 (1)
$10^{4}$ 1.20 (5) 1.20 (1) 1.20 (1)
$10^{-4}$ $10^{-9}$ $\int^{X}$ 1.20 (6) 1.20 (6) 1.20 (6)
1 1.20 (6) 1.20 (6) 1.20 (6)
$10^{4}$ 1.20 (7) 1.20 (7) 1.20 (7)
1 $10^{-9}$ $\int^{X}$ 1.20 (6) 1.21 (6) 1.20 (6)
1 1.20 (7) 1.20 (7) 1.20 (7)
$10^{4}$ 1.98 (19) 1.73 (17) 1.52 (13)
Table 6.4: Example 4. Estimated condition number and MinRes iterations required to converge, associated with the preconditioned system matrix of the linearised Darcy–Forchheimer equations using the preconditioner induced by (5.8).
We stress that the inverse of each block in the preconditioner is applied in an abstract operator sense. In particular, we use the LinearOperator abstraction in Gridap/Julia to define the (linearised) weak form associated with each block of the preconditioner. This allows us to apply a block-diagonal approximation of the preconditioning without explicitly assembling or inverting the full operator. In the numerical examples presented, each block is inverted using the sparse direct solver UMFPACK. We have not employed multigrid solvers, as the numerical examples are of moderate size and are intended to validate the theoretical framework rather than to optimise performance. Nevertheless, the proposed preconditioners are fully compatible with multigrid-based approximations, and we expect the robustness with respect to $\kappa$ and $\mathtt{F}$ to be retained under standard spectral equivalence assumptions.